# Neural Force Field: Learning Generalized Physical Representation from a Few Examples

### v1

@@ 3.2 Interactive planning
Forward planning  For physical reasoning tasks requiring specific action sequences, we extend the NFF framework to incorporate action effects by including action features in $f_{\theta}$. Through forward simulation with the NFF model, we sample multiple action sequences $A=\{a_{0},...,a_{T}\}$ and optimize according to evaluation metrics $R$:
[…]
The optimal sequence $A^{\star}$ can then be executed in the physical environment. While the forward simulation may initially deviate from actual observations, NFF’s physics-based design enables efficient adaptation through new experimental data. This capability for rapid model refinement mirrors human reasoning (Allen et al., 2020) and overcomes the catastrophic forgetting challenges faced by pure neural network approaches (Dohare et al., 2024). When executed trajectories deviate from the desired goal state, the new state sequences can be directly incorporated into model optimization through the MSE loss, following the same training procedure used in the initial learning phase.
Backward planning  By inverting the ODE integrator, NFF enables the computation of initial conditions given a desired goal state. Through backward integration, the initial conditions can be determined analytically:
[…]
The invertible nature of the NFF formulation makes this backward computation particularly efficient. By performing temporal integration in reverse, we obtain initial conditions consistent with both the goal state and the learned physical dynamics encoded in the NFF model.
(a) Support interaction
(b) Platform orientation
(c) Impact dynamics
(d) Spring mechanics
Figure 4: Force response analysis under controlled state variations. Each plot shows the predicted force components (Fx in blue, Fy in orange) from our trained NFF model, measured in normalized units against different state parameters. The subplots systematically analyze: (a) support forces as a function of ball height, showing characteristic contact response when the ball meets the platform, (b) force decomposition as the platform rotates from horizontal ($0^{\degree}$) to vertical ($90^{\degree}$), (c) impact forces scaling with the ball’s downward velocity, and (d) spring forces varying with horizontal displacement from the equilibrium position. Each plot varies one parameter while keeping all other scene variables constant, demonstrating NFF’s learned physical principles, including contact mechanics, geometric reasoning, impact dynamics, and harmonic motion.
@@ 4.1 Environments and datasets
We evaluate our approach on two physical reasoning tasks: I-PHYRE and N-body dynamics. Each task is evaluated across three distinct settings: within-scenario prediction, cross-scenario prediction, and planning. Detailed environment and dataset configurations are provided in Appendix A.
I-PHYRE   I-PHYRE (Li et al., 2023) presents a suite of complex physical reasoning puzzles requiring multi-step interventions. The environment incorporates diverse physical interactions including gravity, collision, friction, rotation, spring dynamics, and pendulum motion. It challenges AI agents to solve puzzles with minimal environmental interactions while generalizing to unseen scenarios. For within-scenario prediction, we evaluate on 10 training games that share similar scenarios but require different solutions. The cross-scenario prediction setting extends to 30 novel games featuring noise, compositional elements, and multi-ball scenarios, with varying object properties such as block lengths, ball sizes, and object positions. In the planning setting, models must generate optimal action sequences to successfully complete each game.
N-body  The N-body problem (Newton, 1833) tests trajectory prediction for small comets orbiting a massive central planet. Using REBOUND simulator (Rein & Liu, 2012), we generate dynamics data by randomly sampling orbital parameters including radii, angles, and masses. This task evaluates the model’s ability to infer gravitational laws from limited observations. The within-scenario prediction setting introduces novel initial conditions and masses, focusing on systems with 1-2 orbiting bodies tracked over 50 timesteps. Cross-scenario prediction significantly increases complexity by introducing systems with 8 orbiting bodies tracked over 100 timesteps. The planning setting challenges models to optimize initial conditions that will evolve to specified target states after 50 timesteps.
@@ 4.2 Learning force fields from a few examples
We qualitatively evaluate NFF’s ability to learn force fields from limited training data; training implementation details are provided in Appendix B. For I-PHYRE, we train NFF on just 10 basic games with 10 trajectory samples each. From these 100 trajectories, the model successfully learns fundamental physical concepts including gravity, support, sliding, collision, friction, and spring dynamics, representing them through unified force fields (Figure 2(g)). To verify that NFF learns generalizable physical principles, we examine force responses to varied ball-block interactions under controlled conditions. Figure 4 shows our systematic evaluation where we independently vary individual scene parameters such as position, velocity, and angle while holding others constant, demonstrating NFF’s accurate force response modeling.
For the N-body system, we train NFF using 200 randomly sampled trajectories from 2-body and 3-body dynamics. As shown in Figure 2(i), the model successfully learns to capture the inverse gravitational field, correctly modeling the distance-dependent centripetal forces governing the mutual attraction between bodies.
(a) I-PHYRE
(b) N-body
Figure 5: NFF outperforms other baselines in prediction tasks. Performance comparison between IN (Battaglia et al., 2016), SlotFormer (Wu et al., 2023), and our NFF on (a) I-PHYRE and (b) N-body tasks. Box plots display median, quartiles, and outliers, with lower values $\downarrow$ better for error metrics (RMSE, FPE, PCE) and higher values $\uparrow$ better for correlation (R). Our NFF consistently achieves lower errors and higher correlations across both simulation environments, particularly in challenging cross-scenario generalization.
@@ 4.3 Prediction on unseen scenarios
We evaluate the learned force fields’ predictive accuracy in both within-scenario and cross-scenario settings. For I-PHYRE, we test against 20 ground truth trajectories per game, while for N-body systems, we evaluate using 200 novel initial conditions. We compare our results against established interaction modeling methods: IN (Battaglia et al., 2016) and SlotFormer (Wu et al., 2023).
Evaluation metrics  We employ multiple complementary metrics to assess prediction quality. Beyond standard Root Mean Squared Error (RMSE), we introduce Final Position Error (FPE) (final position error) and Position Change Error (PCE) (position change error) for detailed performance analysis. FPE quantifies terminal position accuracy, while PCE evaluates the model’s ability to capture motion dynamics. Additionally, Pearson Correlation Coefficient (R) measures trajectory shape alignment independent of speed variations. This comprehensive metric suite enables thorough evaluation across temporal and spatial dimensions. Detailed definitions are provided in Section B.2.
Results  As shown in Figure 3, NFF generates physically plausible trajectories that closely match ground truth behavior, even in previously unseen scenarios. Quantitative comparisons in Figure 5 demonstrate NFF’s superior performance across all metrics for both I-PHYRE and N-body tasks, with particularly strong results in cross-scenario generalization. While SlotFormer exhibits overfitting tendencies that limit its cross-scenario performance (analyzed further in Appendix C), NFF’s ability to learn generalizable physical principles from limited training data enables robust prediction across diverse scenarios.
@@ 4.4 Planning on unseen scenarios
I-PHYRE planning  We implement a 5-round interactive learning protocol. NFF acts as a mental simulator to evaluate 500 randomly sampled action candidates, selecting the optimal sequence for physical execution. After each execution, the model updates its parameters based on observed outcomes, refining its physics understanding. This updated model then guides subsequent action proposals.
Figure 6: Interactive planning performance on I-PHYRE. Comparison of cumulative success probability over 5 planning rounds between different approaches: human performance random sampling, our NFF, IN, and SlotFormer. Solid and dashed lines indicate variants with and without the refinement mechanism (Allen et al., 2020), respectively. Error bars show SEM across trials. Our NFF with refinement (yellow) demonstrates continuous improvement across rounds, achieving performance comparable to human players (data from Li et al. (2023)), while baseline models show limited performance even with refinement.
(a) Euler step size
(b) Step size $25e-$ 2
(c) Step size $35e-$ 3
(d) Latent dimension
(e) Ground-truth
(f) Dim = 3
(g) Dim = 16
(h) Dim = 256
(i) Latent space
Figure 7: Ablation studies on key model components in N-body systems. Results averaged across test cases and five random seeds. (a) Model error vs. integration step size ($31e-$ 3 to $25e-$ 2), showing improved accuracy with finer Euler integration. (b-c) Performance comparison with and without ODE integration for within- and cross-scenario generalization, evaluated at different step sizes. (d) Impact of latent space dimensionality (3, 16, 256) on generalization, demonstrating optimal performance when matching the inherent 3D nature of gravitational forces. (e) UMAP visualization of learned force field representations, revealing that lower dimensionality better captures the underlying gravitational field manifold compared to higher dimensions. Blue/orange indicates relative magnitude of projected interactions.
@@
We quantitatively evaluate planning performance by calculating success probability $p_{i}$ for each game as the success rate over 20 trials in round $i$, with NFF updating after failures; see detailed results in Appendix E. Figure 6 compares NFF against random sampling, human performance (from Li et al. (2023)), IN, and SlotFormer using cumulative success probability: $1-\prod_{i=1}^{n}(1-p_{i})$ for each trial $n$. NFF outperforms baselines and approaches human-level performance after refinement (Allen et al., 2020), demonstrating effective few-shot learning. The poor performance of IN and SlotFormer, falling below random sampling, indicates how inaccurate dynamics modeling compromises planning.
@@
N-body planning  We focus on determining initial conditions that achieve desired final configurations under celestial dynamics. NFF’s inverse simulation capability enables direct computation of initial conditions through reverse time evolution, while IN and SlotFormer rely on iterative gradient-based refinement. As shown in Table 1, NFF achieves superior planning performance in both within-scenario and cross-scenario settings.
@@
Table 1: N-body system initial condition reconstruction from target configurations. MSE $\downarrow$ comparison between IN, SlotFormer, and our NFF. Results show average and SEM across trials for both within- and cross-scenario evaluations. Lower values indicate better reconstruction accuracy.
@@
Within $\downarrow$ 0.651 $\pm$ 0.021 0.837 $\pm$ 0.029 0.067 $\pm$ 0.010
Cross $\downarrow$ 4.654 $\pm$ 0.193 4.018 $\pm$ 0.133 0.140 $\pm$ 0.011
@@ 4.5 Ablation study
We investigate three key factors affecting model performance: integration precision, ODE grounding, and latent dimensionality. For integration precision, we evaluate NFF on the N-body task using Euler integration at four precision levels: $31e-$ 3, $35e-$ 3, $21e-$ 2, and $25e-$ 2. Figure 7(a) demonstrates that higher integration precision consistently yields better performance. We also explore more integration methods in Appendix D. To study ODE grounding’s impact on generalization, we compare two configurations: our NFF using predicted force fields with ODE integration, and IN using learned state transitions at matching step sizes. Figures 7(b) and 7(c) shows that ODE integration enhances generalization performance. However, at coarser step sizes ($25e-$ 2), NFF slightly underperforms IN, indicating that ODE-based approaches require fine integration precision to realize their advantages. For latent dimensionality analysis, we vary the latent interaction space dimension in NFF across 3, 16, and 256 dimensions. Results in Figure 7(d) reveal that lower-dimensional representations improve generalization capability. The visualization in Figure 7(i) confirms this finding, showing that reduced dimensionality leads to latent representations more closely matching ground truth. These ablation studies demonstrate that NFF achieves optimal few-shot learning of physical dynamics when combining high integration precision, low-dimensional latent space, and ODE-based integration.

### v4

@@ 3.2 Interactive planning
Forward planning  For physical reasoning tasks requiring specific action sequences, we extend the NFF framework to incorporate action effects by including action features in $f_{\theta}$. Through forward simulation with the NFF model, we sample multiple action sequences $A=\{a_{0},...,a_{T}\}$ and optimize according to evaluation metrics $R$: $A^{\star}=\underset{A}{\text{argmax}}R(\mathbf{F},A)$. The optimal sequence $A^{\star}$ can then be executed in the physical environment. While the forward simulation may initially deviate from actual observations, NFF’s physics-based design enables efficient adaptation through new experimental data, inspired by human trial-and-error learning (Allen et al., , 2020). When executed trajectories deviate from the desired goal state, the new state sequences can be directly incorporated into model optimization through the MSE loss, following the same training procedure used in the initial learning.
Backward planning  The invertible nature of the NFF formulation makes backward computation particularly efficient. By inverting the time direction of ODE integrator, NFF enables the inversion of initial conditions given a desired goal state: $\mathbf{x}(0)=\mathbf{x}(t)+\int_{t}^{0}\mathbf{v}(t)\,dt$, $\mathbf{v}(0)=\mathbf{v}(t)+\int_{t}^{0}\mathbf{F}(\mathbf{z}^{q}(t))\,dt$, consistent with the learned physical dynamics encoded in the NFF model, as shown in Fig.˜10.
@@ 4.1 Environments and datasets
We evaluate state-based NFF on I-PHYRE and N-body dynamics and vision-based NFF on PHYRE. Each task is evaluated across three distinct settings: within-scenario prediction, cross-scenario prediction, and planning. Detailed environment and dataset configurations are provided in Appx.˜A.
I-PHYRE   I-PHYRE (Li et al., , 2023) presents a suite of complex physical reasoning puzzles requiring multi-step interventions. The environment incorporates diverse physical interactions including gravity, collision, friction, rotation, spring dynamics, and pendulum motion. It challenges AI agents to solve puzzles with minimal environmental interactions while generalizing to unseen scenarios. For within-scenario prediction, we evaluate on 10 training games that share similar scenarios but require different solutions. The cross-scenario prediction setting extends to 30 novel games featuring noise, compositional elements, and multi-ball scenarios, with varying object properties such as block lengths, ball sizes, and object positions. In the planning setting, models must generate optimal action sequences to successfully complete each game.
N-body  The N-body problem (Newton, , 1833) tests trajectory prediction for small comets orbiting a massive central planet. Using REBOUND simulator (Rein and Liu, , 2012), we generate dynamics by randomly sampling orbital parameters including radii, angles, and masses. This task evaluates the model’s ability to infer gravitational laws from limited observations. The within-scenario prediction introduces novel initial conditions and masses, focusing on systems with 1-2 orbiting bodies tracked over 50 timesteps. Cross-scenario prediction significantly increases complexity by introducing systems with 7 or 9 orbiting bodies tracked over 150 timesteps. The planning setting challenges models to optimize initial conditions that will evolve to specified target states after 50 timesteps.
PHYRE  Building on the previous two datasets that used state-based representations of objects, we explore extending NFF to a vision-based physical reasoning benchmark, PHYRE (Bakhtin et al., , 2019). Designed for goal-oriented physical reasoning, PHYRE features 25 task templates involving interactive physics puzzles, all solvable by strategically placing a red ball. Each template generates 100 variations of a core task, supporting two distinct evaluation protocols: (i) the within-scenario setting, where models train on 80% of the tasks in each template and are tested on the remaining 20%; and (ii) the cross-scenario setting, which evaluates generalization by training on all tasks from 20 templates and testing on tasks from the 5 unseen ones.
Ground truth
Learned one
(a) PHYRE force field
(b) Gravitational field
Figure 5: Visualization of learned force fields after few-shot learning. (a) NFF successfully inverts force fields across different templates on PHYRE, consisting of physical behaviors like falling under gravity, sliding with friction, and colliding with momentum transfer. (b) The learned gravitational field distributions in N-body also show accurate force field reconstruction.
True
NFF
IN
SlotFormer
I-PHYRE N-body
Figure 7: Trajectory predictions on unseen scenarios after few-shot learning. After learning from 100 and 200 trajectories on I-PHYRE and N-body systems, respectively, our NFF predictions closely match the ground truth behaviors across diverse scenarios, from rigid body interactions to gravitational dynamics. In contrast, other baselines fail to predict physically plausible dynamics. Additional visualizations are provided in Appx.˜G.
@@ 4.2 Learning force fields from a few examples
We qualitatively evaluate NFF’s ability to inverse force fields from limited observations without direct access to ground-truth forces; training details are provided in Appx.˜B. For I-PHYRE, we train NFF on 10 basic games with 10 trajectory samples each. From these 100 trajectories, the model successfully learns fundamental physical concepts through unified force fields. To verify that NFF learns generalizable physical principles, we examine force responses to varied ball-block interactions under controlled conditions. Fig.˜A1 shows our systematic evaluation where we independently vary individual parameters such as position, velocity, and angle while holding others constant, demonstrating NFF’s accurate force response modeling. For the N-body system, we train NFF using 200 randomly sampled trajectories from 2-body and 3-body dynamics. As shown in Fig.˜5(b), the model successfully learns to capture the inverse gravitational field, correctly modeling the distance-dependent centripetal forces governing the mutual attraction between bodies. For PHYRE, we train a vision-based NFF on 12,00012{,}000 trajectories (20 templates $\times$ 20 tasks $\times$ 30 actions), which is about 267 times less data than prior SOTA (Qi et al., , 2021). As shown in Fig.˜5(a), the model accurately captures force fields across diverse templates.
Table 1: NFF outperforms baselines on prediction tasks. Comparison of NFF, IN (Battaglia et al., , 2016), GCN (Kipf and Welling, , 2016), and SlotFormer (Wu et al., , 2023) on I-PHYRE and N-body tasks. Results are reported with SEM; lower is better ($\downarrow$) for error metrics (RMSE, FPE, PCE), and higher is better ($\uparrow$) for correlation (R). NFF consistently yields lower errors and higher correlations, especially under cross-scenario generalization. The [0,3T] settings extend prediction from 50 to 150 steps.
Metric Model I-PHYRE N-body
Within Cross Within [0,T] Within [0,3T] Cross [0,3T]
RMSE $\downarrow$ IN 0.124 ± 0.007 0.194 ± 0.004 0.306 ± 0.022 0.832 ± 0.034 4.727 ± 0.111
GCN 0.118 ± 0.008 0.213 ± 0.006 0.387 ± 0.028 1.090 ± 0.050 2.130 ± 0.046
Slotformer 0.067 ± 0.006 0.206 ± 0.005 0.257 ± 0.021 1.117 ± 0.042 3.531 ± 0.128
NFF 0.048 ± 0.004 0.131 ± 0.004 0.139 ± 0.012 0.558 ± 0.024 1.090 ± 0.023
FPE $\downarrow$ IN 0.129 ± 0.010 0.178 ± 0.005 0.283 ± 0.019 0.740 ± 0.026 5.376 ± 0.152
GCN 0.118 ± 0.009 0.213 ± 0.006 0.332 ± 0.023 1.028 ± 0.048 2.147 ± 0.046
Slotformer 0.068 ± 0.007 0.198 ± 0.005 0.250 ± 0.018 1.289 ± 0.066 3.708 ± 0.134
NFF 0.042 ± 0.004 0.111 ± 0.004 0.147 ± 0.013 0.506 ± 0.022 1.166 ± 0.024
PCE $\downarrow$ IN 0.004 ± 0.000 0.005 ± 0.000 0.018 ± 0.002 0.034 ± 0.003 0.060 ± 0.002
GCN 0.003 ± 0.000 0.194 ± 0.000 0.021 ± 0.002 0.034 ± 0.002 0.041 ± 0.002
Slotformer 0.002 ± 0.000 0.004 ± 0.000 0.015 ± 0.002 0.039 ± 0.002 0.075 ± 0.002
NFF 0.001 ± 0.000 0.003 ± 0.000 0.008 ± 0.001 0.022 ± 0.002 0.032 ± 0.002
R $\uparrow$ IN 0.993 ± 0.001 0.974 ± 0.001 0.968 ± 0.004 0.870 ± 0.011 0.325 ± 0.020
GCN 0.991 ± 0.001 0.969 ± 0.002 0.956 ± 0.005 0.834 ± 0.013 0.586 ± 0.023
Slotformer 0.997 ± 0.001 0.967 ± 0.001 0.977 ± 0.003 0.870 ± 0.010 0.420 ± 0.024
NFF 0.997 ± 0.001 0.980 ± 0.002 0.992 ± 0.001 0.938 ± 0.007 0.826 ± 0.011
@@ 4.3 Prediction on unseen scenarios
We evaluate the learned force fields’ predictive accuracy in both within- and cross-scenario settings. For I-PHYRE, we test against 20 ground truth trajectories per game, while for N-body, we evaluate using 200 unseen initial conditions. We compare against established interaction modeling methods: IN (Battaglia et al., , 2016), GCN (Kipf and Welling, , 2016), and SlotFormer (Wu et al., , 2023).
Evaluation metrics  We employ multiple complementary metrics to assess prediction quality. Beyond standard Root Mean Squared Error (RMSE), we introduce Final Position Error (FPE) and Position Change Error (PCE) for detailed performance analysis. FPE quantifies terminal position accuracy, while PCE evaluates the model’s ability to capture motion dynamics. Additionally, Pearson Correlation Coefficient (R) measures trajectory shape alignment independent of speed variations. This comprehensive metric suite enables thorough evaluation across temporal and spatial dimensions. Detailed definitions are provided in Sec.˜B.2.
Results  As shown in Fig.˜7, NFF generates physically plausible trajectories that closely match ground truth behavior, even in previously unseen scenarios. Quantitative results in Tab.˜1 demonstrate NFF’s superior performance across all metrics for both I-PHYRE and N-body tasks, with particularly strong results in cross-scenario generalization. While other models exhibits overfitting tendencies that limit the cross-scenario performance, NFF’s ability to learn generalizable physical principles enables robust prediction across diverse scenarios. Additional results in Appx.˜D demonstrate the superior robustness of our model under noisy state conditions.
True
NFF
RPIN
SlotFormer
PHYRE within PHYRE cross
Figure 9: Trajectory Predictions on PHYRE. We compare the vision-based NFF, trained on 12,00012{,}000 trajectories, with SOTA methods trained on $3.2$ million trajectories. Prior methods decoding objects from latent spaces often suffer from object inconsistency. For example, in the last column, RPIN mistakenly transforms a gray cup into a gray ball, indicating overfitting to training shapes. SlotFormer also shows object disappearance in cross-scenario tasks. In contrast, NFF produces more accurate predictions while maintaining object consistency.
@@ 4.4 Planning on unseen scenarios
I-PHYRE planning  We implement a 5-round interactive learning protocol. NFF acts as a mental simulator to evaluate 500 randomly sampled action candidates (when to eliminate which block in order to drop the balls into the abyss), selecting the optimal sequence for physical execution. After each execution, the model updates its parameters based on observed outcomes, refining its physics understanding. This updated model then guides subsequent action proposals.
@@
We quantitatively evaluate planning by calculating success probability $p_{i}$ for each game as the success rate over 20 trials in round $i$, with NFF updating after failures; see detailed results in Appx.˜F. Figure˜10 compares NFF against random sampling, human performance (from Li et al., (2023)), IN, and SlotFormer using cumulative success probability: $1-\prod_{i=1}^{n}(1-p_{i})$ for current round $n$. NFF outperforms baselines and approaches human-level performance after refinement (Allen et al., , 2020), demonstrating effective few-shot learning. The performance of IN and SlotFormer, falling below random sampling, indicates how inaccurate dynamics modeling compromises planning.
Figure 10: Interactive planning performance on I-PHYRE. Comparison of cumulative success probability over 5 planning rounds among human, random sampling, our NFF, IN, and SlotFormer. Error bars show SEM across trials. Our NFF with refinement (Allen et al., , 2020) shows continuous improvement across rounds, achieving performance comparable to human (data from Li et al., (2023)), while others show limited performance even with refinement.
@@
N-body planning  We focus on determining initial conditions that achieve desired final configurations under celestial dynamics. NFF’s inverse simulation enables direct computation of initial conditions through reverse time evolution, while IN and SlotFormer rely on iterative gradient-based refinement. As shown in Tab.˜3, NFF achieves superior planning performance in both within- and cross-scenario settings.
PHYRE planning  In PHYRE, we extend NFF to learn dynamics directly from RGB videos with segmentation masks by applying sparse convolutions to extract geometric features from the masked objects. Following PHYRE’s standard evaluation protocol, we first utilize NFF to predict trajectories and then use a pretrained TimeSFormer (Bertasius et al., , 2021) as the classifier to determine task success or failure. Planning performance is measured by the AUCCESS metric (Bakhtin et al., , 2019), which captures both effectiveness and efficiency by assigning higher weight to solutions that succeed with fewer attempts. Specifically, each attempt count $k\in{1,2,3,\dots,100}$ is weighted by $\omega_{k}=\log(k+1)-\log(k)$, and the final AUCCESS score is computed as: $\frac{\sum_{k}\omega_{k}\cdot s_{k}}{\sum_{k}\omega_{k}}$, where $s_{k}$ is the success rate within $k$ attempts.
We compare NFF against the RPIN (Qi et al., , 2021) and SlotFormer (Wu et al., , 2023). As shown qualitatively in Fig.˜9, NFF, trained on just 0.012 million trajectories, learns physically grounded dynamics that support effective planning while maintaining higher fidelity and stronger object permanence than RPIN and SlotFormer, particularly in cross-scenario generalization. Quantitative results in Tab.˜3 further demonstrate that NFF outperforms state-of-the-art baselines in the challenging cross-scenario tasks. In contrast, SlotFormer’s slot attention leads to limited generalization in unseen objects, resulting in planning failures in cross scenarios; see Fig.˜A8 for more analysis.
@@
Table 2: N-body system initial condition reconstruction from target configurations. Results show average MSE $\downarrow$ with SEM across trials for both within- and cross-scenario evaluations.
@@
Within $\downarrow$ 0.651 ± 0.021 0.837 ± 0.029 0.067 ± 0.010
Cross $\downarrow$ 4.654 ± 0.193 4.018 ± 0.133 0.140 ± 0.011
Table 3: Planning performance on PHYRE in the cross-scenario setting. Training sample size (in millions) and AUCCESS $\uparrow$ score are reported. Our NFF achieves the best performance with the least data.
Metric DQN RPIN SlotFormer NFF
Sample (M) 200.000 3.200 0.120 0.012
AUCCESS 30.96 36.58 21.04 49.22
@@ 4.5 Ablation study
We investigate three key factors that may affect NFF performance, including integration precision, ODE grounding, and NOL. For integration precision, we evaluate NFF on the N-body task using Euler integration at four precision levels: 1e31e-3, 5e35e-3, 1e21e-2, and 5e25e-2. Tab.˜4 demonstrates that higher integration precision consistently yields better performance. We also explore more integration methods in Appx.˜C. To study ODE grounding’s impact on generalization, we compare two configurations: our NFF using predicted force fields with ODE integration, and IN using learned state transitions at matching step sizes. Tab.˜4 shows that ODE integration enhances generalization performance. For NOL, we replace the DeepONet in NFF with a vanilla MLP. The results in the last column of Tab.˜4 show a consistent increase in generalization errors across settings compared with NFF with DeepONet (column one). These ablation studies demonstrate that NFF achieves optimal few-shot learning of physical dynamics when combining high integration precision, ODE-based integration, and NOL.
Table 4: Ablation results emphasize the need for high integration precision, ODE grounding, and NOL. We test NFF with different integration precision. The NFF without ODE grounding degenerates to IN with the same integration precision (1e-3) and the NFF without NOL utilize an MLP as force predictor instead of DeepONet. NFF with high integration precision, ODE, and NOL yields the best generalization results.
Scenario NFF 1e-3 NFF 5e-3 NFF 1e-2 NFF 5e-2 w/o ODE w/o NOL
Training 0.083±0.009 0.137±0.014 0.129±0.016 0.231±0.022 0.044±0.005 0.056±0.005
Within 0.095±0.008 0.139±0.013 0.136±0.014 0.239±0.020 0.354±0.023 0.120±0.021
Within [0,3T] 0.471±0.025 0.558±0.024 0.535±0.027 0.693±0.030 0.949±0.041 0.540±0.042
Cross 1.024±0.021 1.090±0.023 1.084±0.024 1.315±0.033 3.518±0.061 1.347±0.128